find the equation of the line that contains the given point p(-2,1) and has the given slope m=3/2
Answer by lwsshak3(11628)

find the equation of the line that contains the given point p(-2,1) and has the given slope m=3/2
**
standard form of equation for a straight line: y=mx+b, m=slope, b=y-intercept
For given problem:
slope,m=3/2
equation: y=3x/2+b
solve for b using coordinates of given point (-2,1)
1=3*-2/2+b
1=-6/2+b
1=-3+b
b=4
equation: y=3x/2+4